OpenStudy (anonymous):

What is the equation of the parabola, in vertex form, with focus at (2,-4) and directrix y = -6?

OpenStudy (anonymous):

\[(x-2)^{2}=\frac{ 1 }{ 4}(y+5)\]\[(x-2)^{2}=4(y+5)\] \[(x+2)^{2}=\frac{ 1 }{ 4}(y+4)\]\[(y+4)^{2}=-4(x-2)\]

OpenStudy (anonymous):

Is it the first answer?

OpenStudy (anonymous):

oh the first answer is also -1/4